Supplemental Operations Report 20May2023

[title Supplemental Operations Report – May 20th]

Supplemental Operations Report 20May2023

Name of person filing report: Sergii Iakymov

Reason for Report: Routine.

Non-Nominal Systems: Robotic observatory.
Action taken for non-nominal systems: New parts for Robotic observatory are received and installation is pending.

Power system:
Solar: nominal. Input breaker of one of MPTT was off two times this week. After resetting the breaker and solar charger work was nominal until the next bright and hot day. This issue has to be investigated by a certified technician.SOC Last 24 hours: Max 100%; Min 46 %; Avg 71.2%.
VDC Last 24 hours: Max 58.04V; Min 44.34V; Avg 50.17V.Generator run time: not used. Generator hours 3810.3.

Propane Reading Station Tank: 75%
Propane Reading Director Tank: 59%
Propane Reading Intern Tank: 71%
Propane Reading Generator Tank: 72%

Water (Static Tank) – 339 gallons
Water in GreenHab – 165 gallons
Water (Outpost tank) – 324 gallons
Water in Science Dome: 0 gallons
Hab Toilet Tank emptied: n/a

Sojourner rover used: Yes.
Hours: 191.7
Beginning Charge: 100
Ending Charge: 100
Currently Charging: Yes
Notes on Rovers: All are plugged in, charged.
ATV’s Used: (Honda, 350.1, 350.2, 300): None
Reason for use: n/a
Oil Added? No
ATV Fuel Used: 0 Gals
Ethanol Free Gasoline — 0 Gallons
# Hours the ATVs were Used today: 0
Notes on ATV: All parked at the outpost and awaiting to be taken to service.

HabCar used and why, where? Yes. To Hanksville for supplies to pick up.

Crew Car used and why, where? Yes. From Grand Junction by crew.

General notes and comments: Nominal.

Dual Split in Science Dome: off due to warmer weather.

Summary of Internet: Nominal

Summary of suits and radios: Nominal

EVA COMMS: Nominal

Campus wide inspection, if action taken, what and why: nothing to report

Summary of General Operations: All nominal.

Summary of Hab operations: Kitchen outlets are not working, inspection will be done after current crew rotation is over.

Summary of GreenHab Operations: Nominal, heater turned off.

Summary of SciDome Operations: Nominal.

Summary of any Observatory Issues: New parts for Robotic observatory are received and installation is pending.

Summary of RAM Operations: Nothing to report

Summary of Outpost Operations:

Director’s trailer: previously reported leak in the bathroom sink pipe was actually located at the pipe-faucet connection; the connection was adjusted to resolve the issue.
The shed was coated a second time to finish outside panting.
The barrel for the burnable trash has been cleaned from ashes.

Summary of Health and Safety Issues: Nominal.

Questions, Concerns, Supplies needed and Requests: Nothing to report

Copyright © The Mars Society. All rights reserved. | Main Site